A particle is moving along a straight line according to the given equation of motion, where s ft is the directed distance of the particle from the origin at t sec. Find the time when the instantaneous acceleration is zero, and then find the directed distance of the particle from the


a. 𝒔 = (πŸπŸπŸ“)/ (πŸπŸ”π’•+πŸ‘πŸ) βˆ’ 2/5π’•πŸ“ , tβ‰₯𝟎

b. 𝒔= πŸ—π’•πŸ + 𝟐 βˆšπŸπ’•+𝟏, tβ‰₯𝟎


1
Expert's answer
2021-12-05T18:17:18-0500

a)

"s=\frac{125}{16t+32}-\frac{2}{5t^5},t\ge0"

Acceleration ="\frac{d^2s}{dt^2}"

"\frac{ds}{dt}(s)=\frac{2}{t^6}-\frac{2000}{(16t+32)^2}"

"\frac{d^2s}{dt^2}(s)=\frac{64000}{(16t+32)^3}-\frac{12}{t^7}"

Time when Instantaneous acceleration is zero,

"\frac{64000}{(16t+32)^3}-\frac{12}{t^7}=0"

"64000t^7-12(16t+32)^3=0"

"64000t^7-12(4096t^3+24576t^2+49152t+32768)=0"

Solving to get "t\approx1.6839"

Time=1.68 seconds

"s=\frac{125}{16t+32}-\frac{2}{5t^5}"

"=\frac{125}{16(1.68)+32}-\frac{2}{5(1.68)^5}"

=0.055 ft

b)

"s=9t^2+2\sqrt{2t}+1,t\ge0"

"\frac{ds}{dt}=18t+\frac{\sqrt{2}}{\sqrt{t}}"

"\frac{d^2s}{dt^2}=18-\frac{1}{\sqrt{2}t^{\frac{3}{2}}}"

"18-\frac{1}{\sqrt{2}t^{\frac{3}{2}}}=0"

"18({\sqrt{2}t^{\frac{3}{2}}})-1=0"

"t^{\frac{3}{2}}=\frac{1}{25.46}"

"t=0.12seconds"

"s=9t^2+2\sqrt{2t}+1"

"=9(0.12)^2+2\sqrt{2(0.12)}+1"

=0.1296+1.9798

=2.1094 ft